OpenStudy (anonymous):

Which of the following lengths could be the sides of a triangle? 6 cm, 18 cm, 12 cm 15 cm, 22 cm, 8 cm 22 cm, 13 cm, 8 cm 18 cm, 6 cm, 12 cm

OpenStudy (anonymous):

@TheSmartOne @Nnesha could you guys help me please? I think the answer would 18 cm, 6 cm, 12 cm.

TheSmartOne (thesmartone):

The sum of the lengths of any two sides of a triangle is greater than the length of the third side. If you take the three sides of a triangle and add them in pairs, the sum is greater than (not equal to) the third side.

OpenStudy (anonymous):

no its b. since 15, 8<22<15+8

OpenStudy (some.random.cool.kid):

a can be too i think.

OpenStudy (anonymous):

Thank you so much for the link @TheSmartOne so would the answer be C?

TheSmartOne (thesmartone):

13+8=? And is it greater than 22?

OpenStudy (some.random.cool.kid):

its b or A i dont think its c

OpenStudy (anonymous):

on a 12+ 6 equals 18 and that is not greater than18 so it can't be a and c 13+18=21 which is smaller than 22

OpenStudy (anonymous):

13 + 8 = 21, It's smaller than 22.

TheSmartOne (thesmartone):

So C is not correct

OpenStudy (anonymous):

22 + 8 = 30, 30 is bigger than 22.

OpenStudy (anonymous):

So would it be B?

Nnesha (nnesha):

when you add two sides that should be greater than of 3 side example like this